# Florent Meyer    Laurent Guichard    Yann Soullard    Denis Coquenet    Guillaume Gravier    Bertrand Coüasnon

### 04 Mar 2026 v1

@@ 2.2.2 Implicit language modeling
Numerous forms of knowledge injection were proposed to provide NN with information about the structure of the input tokens sequence. To better model the strong correlations between adjacent elements in music sequences, Tian et al. [38] group music-specific tokens before embedding them and sending them to a masked-language-modeling transformer encoder. In a similar manner, text tokens can be grouped based on $2$-grams [29] or a word lexicon [8, 9] to encourage the network to use these related tokens together. Conversely, some work explores the reinforcement of the implicit language modeling of an NN by predicting groups of tokens [5, 37]. Similarly in pure NLP, an $n$-gram masking method enhances the integration of language information into masked language modeling by combining three prediction objectives, namely predicting $n$-gram identities as well as each separate token composing the $n$-gram, and detecting tokens replaced by a generator [44]. Such models require vast amounts of data and training time since they are entirely NN-based, and HTR approaches would require annotated text-image pairs to counteract language shift.
@@ 5 Datasets
As shown in Table 1, we conduct experiments on a total of five handwritten word-level dataset splits, four of which are custom splits of the well-known IAM [24] and RIMES [14] datasets while the latter is an industrial use case on name and surname recognition in forms. The custom splits are provided to exacerbate a language shift and its associated bias, as discussed later on. Additionally, we report the $5$-gram character-level perplexity (PPL) on test sets to assess the strength of the shift between source and target language distributions, where a higher PPL denotes a bigger shift. Third to last and second to last columns use a $5$-gram estimated on the source development set, i.e. the union of the training and validation sets. Last column uses a $5$-gram estimated on the target $n$-gram set, denoted by NG. This target NG set was created to estimate our $n$-grams, ensuring that they are not computed directly on target test sets.
@@ 6.2 Training details
WAN consists in a ten-layer FCN encoder and a small two-layer transformer decoder, down from the $6$ and $8$ decoder layers of TrOCR and the DAN. Each of the ten layers comprises a convolution, ReLU activation, instance normalization [39], $2\times 2$ max pooling and dropout [33] with a probability of 0.1. Each layer’s convolution has a $3\times 3$ kernel, a stride of $1$ and padding in such a way that the output has the same shape as the input. Output channels are increased so that the first three layers output $32$ channels, the next three $64$, the following three $128$ and the final layer $256$. The transformer decoder has an internal dimension $d=256$ with $8$ heads. We employ teacher forcing error (TFE) with probability $0.1$ in the WAN’s transformer decoder during training to simulate wrong previous predictions which might occur at inference time. This demonstrated better generalization abilities in HTR [7]. TFE also impacts NGI, as changing the context generates alternative $n$-gram vectors. Through simulating erroneous predictions, TFE restrains the network’s reliance on both internal language modeling and external LM by reducing its confidence into previous predictions. The WAN is trained with a batch size of 128 and a learning rate of $3\times 10^{-4}$}.

### 10 Mar 2026 v2

@@ 2.2.2 Implicit language modeling
Numerous forms of knowledge injection were proposed to provide NN with information about the structure of the input tokens sequence. To better model the strong correlations between adjacent elements in music sequences, Tian et al. [38] group music-specific tokens before embedding them and sending them to a masked-language-modeling transformer encoder. In a similar manner, text tokens can be grouped based on 22-grams [29] or a word lexicon [8, 9] to encourage the network to use these related tokens together. Conversely, some work explores the reinforcement of the implicit language modeling of an NN by predicting groups of tokens [5, 37]. Similarly in pure NLP, an $n$-gram masking method enhances the integration of language information into masked language modeling by combining three prediction objectives, namely predicting $n$-gram identities as well as each separate token composing the $n$-gram, and detecting tokens replaced by a generator [44]. Such models require vast amounts of data and training time since they are entirely NN-based, and HTR approaches would require annotated text-image pairs to counteract language shift.
@@ 5 Datasets
As shown in Table 1, we conduct experiments on a total of five handwritten word-level dataset splits, four of which are custom splits of the well-known IAM [24] and RIMES [14] datasets while the latter is an industrial use case on name and surname recognition in forms. The custom splits are provided to exacerbate a language shift and its associated bias, as discussed later on. Additionally, we report the 55-gram character-level perplexity (PPL) on test sets to assess the strength of the shift between source and target language distributions, where a higher PPL denotes a bigger shift. Third to last and second to last columns use a 55-gram estimated on the source development set, i.e. the union of the training and validation sets. Last column uses a 55-gram estimated on the target $n$-gram set, denoted by NG. This target NG set was created to estimate our $n$-grams, ensuring that they are not computed directly on target test sets.
@@ 6.2 Training details
WAN consists in a ten-layer FCN encoder and a small two-layer transformer decoder, down from the 66 and 88 decoder layers of TrOCR and the DAN. Each of the ten layers comprises a convolution, ReLU activation, instance normalization [39], $2\times 2$ max pooling and dropout [33] with a probability of 0.1. Each layer’s convolution has a $3\times 3$ kernel, a stride of 11 and padding in such a way that the output has the same shape as the input. Output channels are increased so that the first three layers output $32$ channels, the next three $64$, the following three $128$ and the final layer $256$. The transformer decoder has an internal dimension $d=256$ with 88 heads. We employ teacher forcing error (TFE) with probability $0.1$ in the WAN’s transformer decoder during training to simulate wrong previous predictions which might occur at inference time. This demonstrated better generalization abilities in HTR [7]. TFE also impacts NGI, as changing the context generates alternative $n$-gram vectors. Through simulating erroneous predictions, TFE restrains the network’s reliance on both internal language modeling and external LM by reducing its confidence into previous predictions. The WAN is trained with a batch size of 128 and a learning rate of $3\times 10^{-4}$}.